OpenStudy (anonymous):

a rectangular table is five times as long as it is wide. If the area is 80 ft square, find the length and width of the table?

OpenStudy (anonymous):

let the width be x and length be 5x area=l*w 80=(5x)(x) 80=5x^2 x^2=80/5 x=sqrt16 x=4 therefore,width=4 and length=5(4)=20